OpenStudy (anonymous):

Simplify: (4x − 6) − (5x + 1) Polynomials. I got -x - 5 but thats not one of the choices, please help

OpenStudy (anonymous):

that is because when you distribute the "minus sign" the first step should be \[4x-6-5x-1\]

OpenStudy (anonymous):

why'd the addition sign turn into subtraction???

OpenStudy (anonymous):

i am surprised they didn't give that as one of the choices to trick you you have to subtract each term, so it ought to be \(-x-7\)

OpenStudy (anonymous):

that is the distributive property \[a(b+c)=ab+ac\] \[-(5x+1)=-5x-1\]

OpenStudy (anonymous):

i dont understand how you did the first step. i didnt know you change the signs?

OpenStudy (anonymous):

think \(a(b+c)=ab+ac\) with \(a=-1,b=5x,c=1\)

OpenStudy (anonymous):

lets look at \(5x+6-(3x+1)\) now if \(x=10\) this would be \(56-31\)and if you did that subtraction, you wouldn't add the 1 and the 6, you would subtract them to get \(56-31=25\) it is the same with the variables. if you are going to subtract, you have to subtract everything so \(5x+6-(3x+1)=5x+6-3x-1=2x+5\) just like with the numbers

OpenStudy (anonymous):

okaaaay, so if you are subtraaacting the problem, then you have to change all the signs? :/

OpenStudy (anonymous):

yes the first step is to remove the parentheses using the distributive law, i.e. you have to change the sign of everything in the second parentheses for example \[x^2+3x-4-(2x^2+5x-7)=x^2+2x-4-2x^2-5x+7\]

OpenStudy (anonymous):

typo there, sorry \[x^2+3x-4-(2x^2+5x-7)=x^2+3x-4-2x^2-5x+7\] is better

OpenStudy (anonymous):

i'm sorrry.... I still just don get it.... can you do a simpler problem?